# Hybrid Machine-Learning Particle Identification for the ePIC Proximity-Focusing RICH

### v1

@@ 1 Introduction
The proximity-focusing Ring Imaging Cherenkov (pfRICH) serves as a Particle Identification (PID) detector in the backward region, $3.5\lesssim\eta\lesssim-1.5$, of the upcoming ePIC experiment at the Electron–Ion Collider (EIC). The ability to distinguish final-state hadron production in electron–nucleus collisions is crucial for Semi-Inclusive Deep-Inelastic Scattering (SIDIS) measurements. The pfRICH consists of three main components: an aerogel radiator, a mirror system, and a photon sensor plane.
@@
This implies that for a given momentum measured by the tracking system, the Cherenkov angle $\theta_{c}$ depends on the particle mass. This property allows the pfRICH detector to separate pions, kaons, and protons by comparing the measured angle with the expected values. The pfRICH is expected to provide at least $3\sigma$ separation up to 7 GeV/$c$ [Chatterjee:2024zrn].
@@ 4.4 Feature importance analysis
Notably, several CNN latent features appear among the top 20 most important features, demonstrating that the CNN encoder extracts complementary information from hit patterns that are not fully captured by the tabular beam features alone. This observation validates the hybrid strategy of combining deep learning with gradient-boosted decision trees.

### v2

@@ 1 Introduction
The proximity-focusing Ring Imaging Cherenkov (pfRICH) serves as a Particle Identification (PID) detector in the backward region, $3.5\lesssim\eta\lesssim-1.5$, of the upcoming ePIC experiment at the Electron–Ion Collider (EIC). The ability to distinguish final-state hadron production in electron–nucleus collisions is crucial for Semi-Inclusive Deep-Inelastic Scattering (SIDIS) measurements. The pfRICH consists of three main components: an aerogel radiator, a mirror system, and a photon sensor plane as illustrated in Fig. 1.
@@
This implies that for a given momentum measured by the tracking system, the Cherenkov angle $\theta_{c}$ depends on the particle mass (see Fig. 2). This property allows the pfRICH detector to separate pions, kaons, and protons by comparing the measured angle with the expected values. The pfRICH is expected to provide at least $3\sigma$ separation up to 7 GeV/$c$ [Chatterjee:2024zrn].
@@ 4.4 Feature importance analysis
Notably, several CNN latent features appear among the top 20 most important features, demonstrating that the CNN encoder extracts complementary information from hit patterns that is not fully captured by the tabular beam features alone, thereby validating the hybrid approach. To further assess interpretability, the CNN latent space is diagonalized using a principal-component basis, and the feature-importance hierarchy is found to be stable, with beam kinematics and the reconstructed Cherenkov angle remaining dominant, while only the leading CNN mode contributes at a sub-dominant level.